OpenStudy (anonymous):

24 1/2 Divided By 147 Inches?

OpenStudy (anonymous):

I Need How Many Peices Of 24 1/2 Inches Of Ribbon Can I Make Out Of 147 Inches

OpenStudy (anonymous):

6 pieces

OpenStudy (anonymous):

Thank You

OpenStudy (anonymous):

\[\frac{147}{24+1/2}=\frac{147}{\frac{49}{2}}=147*\frac{2}{49}=6 \]